there are 7 red balls and 8 green balls in a bag what is the probability of pulling out two red balls if the balls are not replaced in the bag?
Answer by stanbon(75887)

there are 7 red balls and 8 green balls in a bag what is the probability of pulling out two red balls if the balls are not replaced in the bag?
----
# of ways to select 2 red balls: 7C2 = 21
# of ways to pick 2 balls randomly: 15C2 = 105
---
P(select 2 red) = 21/105 = 7/35 = 1/5
===================
Cheers,
Stan H.